# Chemotaxis-induced phase separation

### v1

@@ Abstract
Source terms interrupt coarsening. — To account for cell growth and death, we now discuss the limit of weakly broken mass conservation (${\varepsilon\ll 1}$). Here, the average total density $\bar{\rho}_{\mathrm{stat}}$ of the stationary pattern is determined by the balance of overall production and degradation, ${\int_{\Omega}\mathrm{d}x\,s(\tilde{\rho})=0}$ (source balance), not by the initial condition. This condition requires ${0=s(\rho_{\mathrm{hss}})}$ at a density ${\check{\rho}<\rho_{\mathrm{hss}}<\hat{\rho}}$. Moreover, for stability of the elementary (half-mesa/-peak) pattern, the source term $s$ must drive any deviations $\bar{\rho}-\bar{\rho}_{\mathrm{stat}}$ from the average total density that balances production and degradation back to zero. We therefore assume net cell growth at low and net death at high densities as, for example, in logistic growth [cf. Fig. 3(a)]. As mass competition amplifies mass differences, and hence deviations in $\bar{\rho}$, a source term as specified above counteracts the mass-competition instability.
Consider the coalescence mode of a peak pattern [Figs. 2(b), 3(a)], the dominant coarsening pathway at large $\bar{\rho}_{\mathrm{stat}}$ in the mKS model [Fig. 2(d)]. If the peak shifts right by $\delta x$, the right plateau shortens, reducing the total production of cells by approximately ${-\varepsilon s(\rho_{-})\delta x}$ while production increases by the same amount to the left of the peak, as follows from integration of Eq. (2). Given that a peak shift $\delta x$ requires the addition/subtraction of mass $(\hat{\rho}-\rho_{-})\delta x$ to the right/left of the peak, the reduced production right and degradation to the left of the peak shifts the peak by ${(\hat{\rho}-\rho_{-})\partial_{t}^{\mathrm{source}}\delta x\approx-\varepsilon s(\rho_{-})\delta x}$. To this source-induced shift process ${\partial_{t}^{\mathrm{source}}\delta x}$ the competition process must be added. Together, one obtains the total mass-competition rate (for ${\varepsilon\ll 1}$; cf. Refs. [37, 17, 18] for the analogous effect in two-component reaction–diffusion systems)
Here, $\sigma_{\mathrm{D}}^{-}(\Lambda)$ denotes the diffusion-limited part of the growth rate $\sigma^{-}$ that is dominant at large peak distances and masses (see Supplemental Material [24]). Thus, the increased source-induced mass influx to the peak from the left compared to the right counteracts the mass transport through the peak caused by the gradients in $\eta$ [Fig. 3(a)]. Mass competition is suppressed if ${\sigma_{\varepsilon}^{-}(\Lambda)\leq 0}$. Because the mass-competition rate ${\sigma_{\varepsilon}^{-}(\Lambda)}$ typically decreases strongly in $\Lambda$ (exponentially in the mKS model), the mass-competition instability is suppressed above a threshold wavelength $\Lambda_{\mathrm{stop}}$ fulfilling ${\sigma_{\varepsilon}^{-}(\Lambda_{\mathrm{stop}})=0}$ [dotted line in Fig. 3(c)]. Similar formulas hold for the competition mode and mesa patterns.
@@
Plateau splitting at larger wavelengths. — As the source strength or the peak distance is increased further, the source term will increasingly deform the stationary profile compared to the pattern of the mass-conserving system because the production or degradation bends the plateaus inwards toward $\rho_{\mathrm{hss}}$ [Fig. 3(b)]. When the maximum of the low-density plateau (or the minimum of a high-density plateau) enters the region of negative nullcline slope ${\partial_{\rho}\eta^{*}<0}$ at wavelengths ${\Lambda\geq\Lambda_{\mathrm{split}}}$, the plateau loses stability, and a new peak or mesa is nucleated [Fig. 3(b)] [38, 39, 17]. Thus, splitting and interrupted coarsening discussed above steer the pattern toward a stationary, periodic configuration with a wavelength $\Lambda$ fulfilling ${\Lambda_{\mathrm{stop}}<\Lambda<\Lambda_{\mathrm{split}}}$ [Fig. 3(c)].
@@
Successive coalescence and splitting. — The scale separation ${\Lambda_{\mathrm{stop}}\ll\Lambda_{\mathrm{split}}}$ can be lost if the source term (approximately) vanishes at a plateau density $\rho_{\pm}$, as in the mKS model with a logistic growth term ${s(\rho)=\rho(p-\rho)}$. Far below the splitting threshold $\Lambda_{\mathrm{split}}$, coalescence is counteracted by production ${\varepsilon s(\rho_{-})\approx\varepsilon\rho_{-}}$ in the low-density plateau ${\rho_{-}\approx 0}$ [Eq. (5); dotted line in Fig. 3(c)]. In contrast, near the splitting threshold, the plateau deformation inducing splitting is driven by a source term ${\sim\varepsilon\rho\gg\varepsilon\rho_{-}}$. Effectively, the source term appears weak for interrupted coarsening but strong for splitting, causing both thresholds to converge already at small source strength $\varepsilon$. Taking into account that the plateau density grows as the splitting threshold $\Lambda_{\mathrm{split}}$ is approached (see Supplemental Material [24]), the threshold of interrupted coarsening $\Lambda_{\mathrm{stop}}$ bends downward [solid blue line in Fig. 3(c)]. Nonetheless, the difference $\Lambda_{\mathrm{split}}-\Lambda_{\mathrm{stop}}$ decreases strongly with $\varepsilon$, well describing the numerical results in the mKS model [Fig. 3(c)].
@@
Taken together, we provide a unified framework for understanding nonlinear patterns and their dynamics in chemotactic systems by introducing the mass-redistribution potential $\eta$. The gradients of this potential drive the migration of the chemotactic species integrating the counteracting effects of random and chemotactic cell movement. We have identified a geometric construction principle, akin to a Maxwell construction, that exploits flux-balance and reactive area balance. Thereby, our construction has a clear connection to the physics of the underlying chemotaxis system. As equilibrium-like pattern constructions were also obtained for different systems from the broad class of scalar active matter [40, 41, 42, 20, 43, 17, 44, 45, 46, 47], it is an exciting question how these constructions can be used to classify the coarse-grained dynamics of scalar active matter more generally.
Keller–Segel-type drift terms also arise if cells modulate motility based on quorum sensing [48, 49]. If instantaneous and (quasi) local this sensing leads to motility-induced phase separation. Also cross-diffusion gives rise to similar drift terms and phase-separation-like behavior [21]. Our results show more broadly that the coarse-grained dynamics of chemotactic systems resembles the phenomenology of phase separation.
Providing a unified understanding of generalized Keller–Segel models, we expect that our theory will provide valuable insights into multi-species chemotactic systems [26, 50, 51], the interplay with different aggregation effects like phase separation [52, 53, 54], and systems with more complex signaling [55]. In particular, it remains an intriguing question whether traveling fronts observed in systems with chemoattractant consumption, rather than production, or nutrient uptake can be analyzed using a similar framework [56, 57, 58, 59]. Moreover, understanding the effect of noise on wavelength selection and the spatiotemporally chaotic steady-state will be an exciting future challenge.

### v2

@@ Abstract
Source terms interrupt coarsening. — To account for cell growth and death, we now discuss the limit of weakly broken mass conservation ($\varepsilon\ll 1$}). Here, the average total density $\bar{\rho}_{\mathrm{stat}}$ of the stationary pattern is determined by the balance of overall production and degradation, $\int_{\Omega}\mathrm{d}x\,s(\tilde{\rho})=0$} (source balance), not by the initial condition. Thus, a pattern only forms if $0=s(\rho_{\mathrm{hss}})$} holds at a density $^{\check{\rho}<\rho_{\mathrm{hss}}<\hat{\rho}}$. Otherwise the system will be homogeneously filled or depleted of cells due to excessive cell growth or death. Moreover, for stability of the elementary (half-mesa/-peak) pattern, the source term $s$ must drive any deviations $\bar{\rho}-\bar{\rho}_{\mathrm{stat}}$ from the average total density that balances production and degradation back to zero. We therefore assume net cell growth at low and net death at high densities as, for example, in logistic growth [cf. Fig. 3(a)]. As mass competition amplifies mass differences, and hence differences in $\bar{\rho}$ between elementary patterns, the source term counteracts the mass-competition instability.
To illustrate the counteracting effect, consider the coalescence of a peak with its neighbor (mirror image) [Figs. 2(b), 3(a)], the dominant coarsening pathway at large $\bar{\rho}_{\mathrm{stat}}$ in the mKS model [Fig. 2(c,d)]. If the peak shifts right by $\delta x$, the right plateau shortens, reducing the total production of cells in the right plateau by approximately $\varepsilon s(\rho_{-})\delta x$} while production increases by the same amount to the left of the peak, as follows from integration of Eq. (2). As a result, the chemotactic flux of new cells into the aggregate, that balances the degradation of cells inside, is asymmetric and larger on the left side of the aggregate. Thus, the aggregate will shift back towards the symmetric position because more cells accumulate on its left side [Fig. 3(a)]. To this stabilizing, source-induced shift process the destabilizing mass-competition process due to the $\eta$ gradients in the first term of Eq. (2) [Fig. 2(b)] must be added. To lowest order in $\varepsilon$, the competition effect can be described using the result for the mass-conserving system (see Appendix D). Mass competition is suppressed if the growth rate of the mass-competition instability under the influence of weak source terms fulfills $\sigma_{\varepsilon}^{-}(\Lambda)\leq 0$}. Because the mass-competition rate $\sigma^{-}(\Lambda)$} typically decreases strongly in $\Lambda$ (exponentially in the mKS model), the mass-competition instability is suppressed above a threshold wavelength $\Lambda_{\mathrm{stop}}$ fulfilling $\sigma_{\varepsilon}^{-}(\Lambda_{\mathrm{stop}})=0$} [dotted line in Fig. 3(c)]. Similar formulas follow from the same reasoning for the competition between peaks and for mesa patterns.
@@
Plateau splitting at larger wavelengths. — As the source strength or the peak distance is increased further, the source term will increasingly deform the stationary profile compared to the pattern of the mass-conserving system. The weak source terms create weak spatial gradients. These gradients let long plateaus bend inward significantly toward $\rho_{\mathrm{hss}}$ [Fig. 3(b), see  [27]]. When the maximum of the low-density plateau (or the minimum of a high-density plateau) enters the region of negative NC slope $\partial_{\rho}\eta^{*}<0$} at wavelengths $\Lambda\geq\Lambda_{\mathrm{split}}$}, the plateau loses stability, and a new peak or mesa is nucleated [Fig. 3(b)] [45, 41, 20]. Thus, splitting and interrupted coarsening steer the pattern toward a stationary, periodic configuration with a wavelength $\Lambda$ fulfilling $\Lambda_{\mathrm{stop}}<\Lambda<\Lambda_{\mathrm{split}}$} [Fig. 3(c)].
@@
Successive coalescence and splitting. — The scale separation $\Lambda_{\mathrm{stop}}\ll\Lambda_{\mathrm{split}}$} can be lost if the source term (approximately) vanishes at a plateau density $\rho_{\pm}$, as in the mKS model with a logistic growth term $s(\rho)=\rho(p-\rho)$}. Far below the splitting threshold $\Lambda_{\mathrm{split}}$, coalescence is counteracted by production $\varepsilon s(\rho_{-})\approx\varepsilon\rho_{-}$} in the low-density plateau $\rho_{-}\approx 0$} [Eq. (5); dotted line in Fig. 3(c)]. In contrast, near the splitting threshold, the plateau deformation inducing splitting is driven by a source term $\sim\varepsilon\rho\gg\varepsilon\rho_{-}$} because the plateau bends upward. Effectively, the source term appears weak for interrupted coarsening but strong for splitting, causing both thresholds to converge already at small source strength $\varepsilon$. Taking into account that the plateau density grows as the splitting threshold $\Lambda_{\mathrm{split}}$ is approached (see [27]), the threshold of interrupted coarsening $\Lambda_{\mathrm{stop}}$ bends downward [solid blue line in Fig. 3(c)]. Nonetheless, the difference $\Lambda_{\mathrm{split}}-\Lambda_{\mathrm{stop}}$} decreases strongly with $\varepsilon$, well describing the numerical results in the mKS model [Fig. 3(c)].
@@
Taken together, we have studied the generic dynamics of chemotactic aggregates using the mass-redistribution potential $\eta$, whose gradients fully determine cell migration by integrating random and chemotactic movement. Our geometric construction principle, akin to a Maxwell construction, leverages flux-balance and reactive area balance, directly linking to the underlying chemotactic physics.
The construction and macroscopic dynamics closely resemble those in two-component mass-conserving reaction–diffusion (2cMcRD) systems [22, 46, 31, 23, 20, 21], although pattern formation in KS models is driven by a nonlinear cross-diffusion term, producing a mass-redistribution potential that depends highly nonlinearly on the cell and chemoattractant densities. We expect that the phenomena likewise extend to higher-dimensional systems [10, 12, 47]. As equilibrium-like pattern constructions were also obtained for scalar active matter [48, 49, 50, 51, 52, 53, 54, 55], it is an exciting question how these constructions can be used to classify the coarse-grained dynamics of scalar non-equilibrium matter more generally.
We expect that our theory will offer valuable insights into multi-species chemotactic systems [29, 56, 57], their interplay with other aggregation mechanisms such as phase separation [58, 59, 60], and systems with more complex signaling [61]. In particular, interrupted coarsening and domain splitting may shape the thymus medulla [57]. The predicted phenomena may also be experimentally accessible using engineered bacteria (see, e.g., Ref. [62]). An intriguing open question is whether chemorepelling systems [63] and traveling fronts in systems with chemoattractant consumption or nutrient uptake can be analyzed using a similar framework [64, 65, 66, 67]. Finally, understanding how noise affects wavelength selection and spatiotemporally chaotic steady states remains an important challenge for future work.
@@ Acknowledgements.
Appendix A Nullcline-slope criterion exact for short-wavelength perturbations
The nullcline-slope criterion predicts the stability of the hss against chemotactic aggregation. It is derived above using a QSS approximation in the long-wavelength limit. For short-wavelength perturbations resulting in fast mass redistribution, the QSS approximation does not include the timescale of the local reactive relaxation of $\eta$ onto $\eta^{*}(\rho)$. Nevertheless, $\eta^{*}(\rho)$ still sets the relaxation direction, and the nullcline-slope criterion, Eq.(4), remains valid: it determines whether $\eta$ increases or decreases in regions of increased cell density $\rho$, and thus whether uniform densities are destabilized [40].
Appendix B Quantifying coarsening of chemotactic aggregates
To understand the dynamics of patterns with several domains (peaks or mesas), we describe the interaction between two elementary pattern motifs (two half-peak or two half-mesa patterns). As illustrated for peaks in the following, the pattern domains can interact either via competition for mass or coalescence with neighbors (mirror image) [Fig. 2(a,b); reflective boundaries]. Assuming a slow evolution of the pattern after each single elementary pattern relaxed to its (quasi-)stationary profile, we approximate the mass-redistribution potential at the single interfaces by its stationary value $\eta_{\mathrm{stat}}$ (QSS approximation). Via the reactive area balance, $\eta_{\mathrm{stat}}$ depends on the extremal densities $\check{\rho}$ and $\hat{\rho}$ of the elementary pattern [Fig. 2(a,b)]. Consequently, differences in peak height or in the width of the upper and lower plateau regions generate $\eta$ gradients: Changes in the plateau lengths shift $\check{\rho}$ or $\hat{\rho}$ exponentially, thereby altering $\eta_{\mathrm{stat}}$. For peak patterns, where the upper plateau is absent, the peak mass directly sets the height $\hat{\rho}$ and thus determines $\eta_{\mathrm{stat}}$. All such changes of the pattern reflect effective changes of the mass $M$ of the elementary pattern, since changes in plateau length due to a shift of the peak or mesa interface [cf. Fig. 2(b)] require mass changes $\delta M\approx\pm(\hat{\rho}-\check{\rho})\,\delta l$}; the sign distinguishes whether the upper or lower plateau is varied. The corresponding partial derivatives of the stationary mass-redistribution potential are denoted by $\partial_{M}^{\pm}\eta_{\mathrm{stat}}$. Since the geometry of the reactive area balance is model-independent, it generically implies $\partial_{\hat{\rho},\check{\rho}}\eta_{\mathrm{stat}}\lessgtr 0$} and, with the above definition of $\delta M$, $\partial_{M}^{\pm}\eta_{\mathrm{stat}}<0$} for both peak and mesa patterns, assuming that $\hat{\rho}$ increases with mass in peaks [Fig. 2(a,b)].
In peak competition [Fig. 2(a)], the mass difference $\delta M$ of the two peaks generates an $\eta$ gradient $\sim\partial_{M}^{+}\eta_{\mathrm{stat}}/\Lambda$} that drives mass redistribution; the larger peak grows, the smaller collapses. In coalescence [Fig. 2(b)], unequal plateau widths shift $\eta_{\mathrm{stat}}$, inducing a gradient $\sim\partial_{M}^{-}\eta_{\mathrm{stat}}/\ell_{\mathrm{int}}$} within the peak of half-width $\ell_{\mathrm{int}}$. This gradient drives peak motion until the shorter plateau vanishes. In both cases, the symmetric configuration is destabilized by a mass-competition instability. Its growth rates $\sigma_{\pm}$ are derived in the Supplemental Material [27] within the QSS approximation and in Ref. [40] using singular perturbation theory. Singular perturbation theory shows that the instability (coarsening) criteria $\partial_{M}^{\pm}\eta_{\mathrm{stat}}<0$ remain exact beyond the QSS approximation.
In terms of diffusion and chemotaxis, what causes generic aggregate coarsening analogously to Ostwald ripening? Linear stability of the low-density plateau requires $\partial_{\rho}\eta^{*}|_{\rho_{-}}>0$} and $\partial_{c}f(\rho_{-},c_{-})<0$}, which together imply $\partial_{\rho}c^{*}|_{\rho_{-}}=-\partial_{\rho}f/\partial_{c}f|_{\rho_{-}}>0$}. Accordingly, $\partial_{M}^{+}\eta_{\mathrm{stat}}<0$} implies $\partial_{M}^{+}\rho_{-}<0$}, such that cells diffuse toward larger aggregates because these deplete the surrounding low-density plateau more strongly. In contrast, chemotaxis drives cells toward smaller aggregates since $\partial_{M}^{+}c_{-}<0$}, but the generic finding $\partial_{M}^{\pm}\eta_{\mathrm{stat}}<0$} implies that the diffusive flux dominates. Taken together, although chemotaxis biases motion of cells towards smaller aggregates, their random motion dominates in the surrounding low-density plateaus and leads to the growth of larger aggregates. These deplete their surroundings more strongly, resulting in aggregate coarsening.
Appendix C Coarsening in the mKS model
In the mKS model, we find $\sigma^{\pm}\sim\exp(-\Lambda/\gamma_{\pm})$} where $\Lambda$ is the wavelength of the pattern [Fig. 2(a,b)] and $\gamma_{\pm}$ are length scales related to the exponential approach of the pattern profile towards $\rho_{\pm}$ (see [27] and Ref. [40]). One has $\gamma_{-}\gtrless\gamma_{+}$} for $\bar{\rho}\gtrless 2D_{\rho}/T$} [Fig. 2(d), inset]. Consequently, coalescence is faster for large average densities $\bar{\rho}>2D_{\rho}/T$} (competition at low densities $\bar{\rho}<2D_{\rho}/T$}), and dominates the coarsening process at large wavelengths. This prediction agrees with the observations in numerical kymographs [Fig. 2(c)]. As $\sigma^{\pm}(\langle\Lambda\rangle)]^{-1}$} is the typical timescale on which peaks vanish via competition or coalescence at the mean pattern wavelength $\langle\Lambda\rangle$ (and strongly increases with $\langle\Lambda\rangle$), the time during coarsening scales as $t^{-1}\sim\sigma^{\pm}(\langle\Lambda\rangle)$} [68, 69, 21]; see [27]. The resulting coarsening law $\langle\Lambda\rangle\sim\max(\gamma_{\pm})\log t$} describes numerical simulations well, with deviations at the transition from competition- to coalescence-dominated coarsening [Fig. 2(d)].
Appendix D Mass competition under the influence of weak source terms
The effect of the source term onto the coalescence of neighboring peaks [Fig. 3(a)] can be analyzed by making a traveling-wave ansatz $\rho(x,t)=\tilde{\rho}(x-\delta x(t))$} for the aggregate. Integrating Eq. (2) from the peak maximum to the right boundary, one finds $(\hat{\rho}-\rho_{-})\partial_{t}^{\mathrm{source}}\delta x\approx-\varepsilon s(\rho_{-})\delta x$}, which drives the peak back to its symmetric position $\delta x=0$} (see [27]). Moreover, to lowest order in $\varepsilon$, the destabilizing mass-competition effect due to $\eta$ gradients within the peak is described by the result $\partial_{t}^{\mathrm{comp}}\delta x\approx\sigma^{-}_{\mathrm{D}}\delta x$ for the mass-conserving system. Together, one obtains the growth rate of the mass-competition instability under the influence of weak source terms (for $\varepsilon\ll 1$}; cf. Refs. [8, 41, 42] and [70, 20, 21] for the analogous effect in two-component reaction–diffusion systems)
Here, $\sigma_{\mathrm{D}}^{-}(\Lambda)$ denotes the diffusion-limited part of the growth rate $\sigma^{-}$. A detailed derivation is given in the Supplemental Material [27].